### v4

# ToddlerBot: Open-Source ML-Compatible Humanoid Platform for Loco-Manipulation

Abstract
Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection. We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for robotics and AI research. ToddlerBot enables seamless acquisition of high-quality simulation and real-world data. The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin and zero-shot sim-to-real policy transfer. A user-friendly teleoperation interface streamlines real-world data collection from human demonstrations. With its data collection ability and anthropomorphic design, ToddlerBot is ideal for whole-body loco-manipulation research. Additionally, ToddlerBot’s compact size ($0.56\penalty 10000\ \mathrm{m},3.4\penalty 10000\ \mathrm{kg}$) ensures safe operation in real-world environments. Reproducibility is achieved with entirely 3D-printed, open-source design and off-the-shelf components, keeping the total cost under $6,000\penalty 10000\ \mathrm{USD}$. This allows assembly and maintenance with basic technical expertise, as validated by successful independent replications of the system. We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, loco-manipulation tasks, and a collaborative long-horizon scenario where two robots tidy a toy session together. By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust and scalable platform for policy learning and execution in robotics research.
Keywords: Humanoid, Mechanisms & Design, Robot Modeling & Simulation
Figure 1: ToddlerBot is an open-source humanoid platform for large-scale, high-quality data collection in both simulation and the real world. It combines massive parallel simulation and an accurate digital twin for simulation, and an intuitive teleoperation device for whole-body control in the real world. ToddlerBot enables diverse loco-manipulation skills, including walking, push-ups, pull-ups, wagon pushing, bimanual, and full-body manipulation, learned from both data sources.
1 Introduction
Conventional robot design prioritizes actuator strength, sensor accuracy, and mechanical precision. While these characteristics are essential for deploying robots in real-world environments, they fall short in addressing the needs of robotics research—where affordability, rapid repairability, and full-stack ownership with no black boxes are crucial for faster innovation, easier debugging, and total system understanding. Additionally, traditional robot platforms are often not aligned with modern machine learning (ML) paradigms driven by embodied data. A robot platform compatible with ML must possess the innate ability to collect observation and action data, both in simulation and the real world, as these complementary data sources are essential for scalable policy learning.
In response to these challenges, we introduce ToddlerBot, an open-source humanoid robot platform for the robotics and AI research community (Figure 1), specifically developed to facilitate policy learning for locomotion and manipulation skills. ToddlerBot is designed to resemble a miniature human to minimize construction and maintenance costs and specialized expertise, lowering the barrier for researchers seeking to build self-sufficient, fully transparent humanoids free from critical reliance on third-party components. Although ToddlerBot lacks the payload capacity of full-sized humanoids, it offers more degrees of freedom in both the upper and lower body than any existing miniature humanoids—sufficient for advancing research in manipulation and locomotion.
Inspired by recent quadrupeds [1, 2] and robotic manipulators [3, 4, 5, 6, 7], ToddlerBot is designed with ML-compatibility in mind. Researchers can use ToddlerBot to not only execute policies but also to serve as a robust data collection platform. It enables the acquisition of high-quality simulation data through a plug-and-play zero-point calibration procedure and transferable motor system identification (sysID) results. These tools ensure a high-fidelity digital twin. We validate the quality of simulation data with keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking), demonstrating the capability for zero-shot sim-to-real transfer. ToddlerBot can also acquire scalable real-world data. We design an intuitive teleoperation interface that allows simultaneous control of ToddlerBot’s upper and lower body to collect whole-body manipulation data and develop effective visuomotor policies. Additionally, ToddlerBot’s small size and weight ($0.56\penalty 10000\ \mathrm{m},3.4\penalty 10000\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
Beyond ML-compatibility, ToddlerBot is designed with a focus on capability and reproducibility. A humanoid has the potential to utilize large scale human demonstrations by leveraging its anatomical similarity to the human body. As such, ToddlerBot features an anthropomorphic design with 30 active degrees of freedom (excluding end effectors). We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, and various open-loop and closed-loop loco-manipulation tasks.
Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components. ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\penalty 10000\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers). We release digital twin, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair. To validate reproducibility, we enlisted a CS-major student not involved in this project to independently build another instance of ToddlerBot. and successfully zero-shot transfer loco-manipulation policies between the two instances. We also received five reports of independent replications of ToddlerBot within a week by teams worldwide.
We summarize our contributions as follows: (1) As a system paper, we integrate existing components to create the first miniature humanoid with 30 DoFs and a superhuman range of motion, offering a practical humanoid platform with new capabilities for the community. (2) We develop a comprehensive sysID pipeline to ensure a high-fidelity digital twin in simulation. (3) We design a whole-body teleoperation solution to collect loco-manipulation data in the real world.
2 Related Works
Table 1: Comparison with Other Popular Humanoid Research Platforms.
Humanoid Size Weight Compute Active DoFs(b) Manipulation Locomotion Sim Data Real Data Open Source Price
(m) (kg) (TFLOPS)(a) - - - - - - (\$)
BD Atlas [8] 1.50 89.0 - 28 - - -
Figure [9] 1.68 70.0 - 26 - - -
Optimus [10] 1.73 57.0 - 28 - - -
Digit [11] 1.75 65.0 - 16 code 250K
Unitree H1 [12] 1.76 47.0 1.92 19 code 70K
Fourier GR1 [13] 1.65 55.0 2.23 32 - - 110K
Booster T1 [14] 1.18 30.0 3.33 23 code 34K
Unitree G1 [15] 1.32 35.0 2.50 29 code 57K
iCub [16] 1.04 24.0 1.93 32 code 300K
Berkeley [17] 0.85 16.0 1.92 12 code 10K
MIT [18] 1.04 24.0 - 18 - - -
Duke [19] 1.00 30.0 Optional 10 design, code 16K
Berkeley Lite [20] 0.80 16.0 0.29 22 design, code 5K
BRUCE [21] 0.70 4.8 0.1 16 code 6.5K
NAO H25 [22] 0.57 5.2 0.02 23 code 14K
Robotis OP3 [23] 0.51 3.5 0.30 20 code 11K
Zeroth [24] 0.48 3.6 0.01 16 design, code 1.4K
Ours 0.56 3.4 2.50 30 design, code 6K
Average Adult [25] 1.73 70.9 - 32(c) - - - -
  • (a)
    TFLOPS refers to Tera FLoating-point Operations Per Second with single precision (FP32).
  • (b)
    The active degrees of freedom actuated by motors, excluding end effectors such as parallel grippers or dexterous hands.
  • (c)
    While human body is powered by over 600 muscles, the primary functional movements can be approximated using 32 revolute joints: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
In recent years, numerous humanoid robots have been developed. Full size industrial humanoids such as Boston Dynamics Atlas [8], Cassie [26], Digit [11], Figure [9], Tesla Optimus [10], Unitree H1 [12] and Fourier GR1 [13]. Half-scaled Booster T1 [14], Berkeley Humanoid Lite [20] and Unitree G1 [15]. Miniature-sized NAO H25 [22], Robotis OP3 [23], and K-Scale Zeroth [24]. Together, they showcase diverse designs and capabilities. On the other hand, humanoids from research institutions, including the Berkeley Humanoid [17], BRUCE [21], Duke Humanoid [19], iCub [16], and MIT Humanoid [18], also explore the humanoid design space with different optimization emphasis. As outlined in Table 1, we identify nine metrics to compare different humanoid platforms.
Size and weight are critical when designing humanoid robots. For manipulation, full-size humanoids are compatible with human-scale objects, whereas reduced-size robots are limited in their ability to manipulate them. However, reduced-size humanoids can still benefit manipulation research when paired with appropriately scaled objects [20]. Furthermore, dynamic whole-body control techniques are likely transferable to larger humanoids. For locomotion, full-sized and small-scale humanoids offer similar research values [27]. Full-size humanoids typically require a substantial engineering team to operate and maintain, along with specialized facilities like gantry cranes for safety. In contrast, a smaller humanoid is inherently cheaper, easier to build and repair, and safer [28, 27, 29]. They can be deployed by a single person and operated in constrained environments with a laptop.
The number of active DoFs is crucial for human-like motion. While the human musculoskeletal system employs over 600 muscles, primary functional movement can be approximated using 32 revolute joints [16, 30]: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes. Therefore, humanoid designs aim to achieve a DoF count as close to 32 as possible. The public perception of the limited performance of miniature humanoids is primarily due to fewer DoFs. This limitation often arises from space constraints that restrict the incorporation of many DoFs, a challenge we have successfully addressed in ToddlerBot.
To qualitatively assess the capability, we evaluate the humanoid’s ability to perform both manipulation and locomotion tasks. The combination of both is particularly compelling, as it unlocks opportunities for whole-body control research [31, 32, 33, 34, 35]. Furthermore, certain motions, such as push-ups, pull-ups, and cartwheels, require coordinated use of both arms and legs. Apart from physical capabilities, ToddlerBot also has best compute in class. Onboard CUDA accelerator allows concurrent policy inferences to support visual and locomotion capabilities. Large-scale simulation data collection is effective for locomotion [36, 37, 38], while real-world data collection is more promising for manipulation [39, 40, 41]. Therefore, to be ML-compatible, a humanoid research platform should facilitate data collection in both world.
Moreover, being open-source and low-cost is essential for others to reproduce. Without these qualities, research in this field would remain restricted to those with specialized expertise and significant resources. While making no compromise in functionality, ToddlerBot stands out as completely open-source and the most accessible among recent humanoid research platforms.
3 System Design
The design space of humanoid robots is vast; ToddlerBot’s key design principles prioritize reproducibility as a hard constraint, and capability and ML-compatibility as key design objectives.
3.1 Reproducibility - A Hard Constraint
Unlike most prior works, we treat reproducibility as a hard constraint, as our humanoid platform holds no value if others cannot reproduce it. We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment. 3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [2, 5, 4]. Despite these merits, it’s challenging to ensure sufficient strength and precision for printed plastic parts. We discuss our justifications for 3D printing in Appendix 8.1. For performance-critical components like motors and bearings, we limit ourselves to off-the-shelf parts. The total BOM cost of ToddlerBot is $6,000\penalty 10000\ \mathrm{USD}$, with 90% spent on the computer (NVIDIA Jetson) and motors (Robotis Dynamixel). We detail our design considerations in Appendix 8.2.
Figure 2: Mechatronic Design. Orange markers highlights ToddlerBot’s 30 active DoFs: 7 per arm, 6 per leg, a 2 on neck, and a 2 on waist. Green markers indicate two end-effector designs—a compliant palm and a parallel-jaw gripper. Purple markers denote the electronics layout with exploded views, featuring 2 fisheye cameras, 1 speaker, 2 microphones, 1 IMU, and 1 Jetson Orin computer.
3.2 Capability - Design Objective
Mechatronic Design. As shown in Figure 2, ToddlerBot replicates the structure of an adult human body to maximize the number of active DoFs, enabling human-like motion and versatile loco-manipulation tasks. Each arm features seven DoFs with spur gears for axis-aligned transmission, ensuring high functionality and reachability. Each leg has six DoFs: three at the hip for a wide range of motion and optimized walking strides, a parallel linkage at the knee to reduce inertia, and two at the ankle for stable locomotion. The neck, with two DoFs, incorporates parallel linkages at the pitch joint for a compact design, allowing expressive motion and full head mobility. The waist, also with two DoFs, uses coupled bevel gears to balance the space budget and effectively transmit power from two motors for yaw and roll actuation, which enables whole-body control. Additionally, ToddlerBot offers two end-effector designs: a parallel jaw gripper for grasping and a compliant palm for tasks requiring a palm-like posture. The end-effector designs can be switched quickly within two minutes. For each active DoF, we maximize the range of motion by optimizing geometries to prevent self-collisions (Appendix 8.3). To address concerns such as space constraints, axis alignment, and inertia reduction, we integrated a combination of three transmission mechanisms in the design: spur gears, coupled bevel gears, and parallel linkages (Appendix 8.4).
Sensors, Compute, and Power. To enhance ToddlerBot’s capabilities, we integrate a comprehensive set of sensors and computational components. Two fisheye cameras are included to expand the field of view. An Inertial Measurement Unit (IMU) is placed in the chest to provide state feedback. A speaker and microphones facilitate communication with humans and other robots. We provide a conversation example between two ToddlerBot instances in the supplementary video. The onboard computation is powered by a Jetson Orin NX 16GB, enabling real-time inference of ML models. Based on our proposed power metric (Appendix 8.5), we choose Dynamixel motors for their reliability and accessibility (Appendix 8.6). Power management is handled by a custom-designed power distribution board which can be easily ordered online (Appendix 8.7).
3.3 ML-Compatibility - Design Objective
Digital Twin. A high-fidelity digital twin is fundamental for high-quality simulation data collection and zero-shot sim-to-real transfer. We divide the digital twin development into two key components: zero-point calibration for correct kinematics and motor system identification for accurate dynamics. Since Dynamixel motors lack an absolute zero point, a reliable method is needed to calibrate after reassembly. We designed a set of calibration devices that can align the robot to the desired zero point within a minute, defined as standing with both arms parallel to the body (Appendix 8.8). For accurate sysID, we designed a motor test bed and automatic test procedures described in Appendix 8.9 to measure the range of dynamics parameters. Next, we collect actuator position tracking data by commanding a chirp signal [27], and use the position tracking result to fit an actuation model [42] with optimization, as detailed in Appendix 8.10. Testing across instances shows Dynamixel motors of the same model have nearly identical dynamics parameters. This is also validated by successful direct deployment of policies to a second ToddlerBot instance without additional sysID.
Teleoperation Device. Inspired by previous works [4, 43], we develop a second upper body of ToddlerBot as the leader arms to collect high-quality real-world data. Two force-sensitive resistors (FSRs) are embedded in the gripping area of the end effectors to detect compression force from the operator, allowing gripper movement based on force input. We use a handheld gaming computer (either Steam Deck or ROG Ally X) to control the other body parts. The joysticks send velocity commands to walk, turn, and squat. Buttons trigger either programmed or trained policies and provide direct control over neck and waist movements. The detailed mapping is described in Appendix 8.11.
4 System Control
4.1 Keyframe Animation
Keyframe animation is widely used in character animation, but it provides only kinematics data, lacking a guarantee of dynamic feasibility [44]. To address this, we develop a tool integrating MuJoCo [45] with a GUI, enabling real-time tuning and validation of keyframe motion trajectories. Combined with high-fidelity digital twin, we can efficiently generate open-loop trajectories such as cuddling, push-ups, and pull-ups that can be executed zero-shot in the real world.
4.2 Reinforcement Learning
For walking, we train a reinforcement learning (RL) policy, $\pi(\bm{\mathrm{a}}_{t}|\bm{\mathrm{s}}_{t})$, which outputs $\bm{\mathrm{a}}_{t}$ as joint position setpoints for proportional-derivative (PD) controllers, based on the observable state $\bm{\mathrm{s}}_{t}$:
$\bm{\mathrm{s}}_{t}=\left(\bm{\phi}_{t},\bm{\mathrm{c}}_{t},\Delta\bm{\mathrm{q}}_{t},\bm{\dot{\mathrm{q}}}_{t},\bm{\mathrm{a}}_{t-1},\bm{\theta}_{t},\bm{\omega}_{t}\right),$ (1)
where $\bm{\phi}_{t}$ is a phase signal, $\bm{c}_{t}$ represents velocity commands, $\Delta\bm{q}_{t}$ denotes the position offset relative to the neutral pose $\bm{q}_{0}$, $\bm{a}_{t-1}$ is the action from the previous time step, $\bm{\theta}_{t}$ represents the torso orientation, and $\bm{\omega}_{t}$ is the torso’s angular velocity. During PPO policy training [46], the environment generates the next state, $\bm{s}_{t+1}$, updates the phase signal, and returns a scalar reward $\mathrm{r}_{t}=\mathrm{r}(\bm{s}_{t},\bm{a}_{t},\bm{s}_{t+1},\bm{\phi}_{t},\bm{c}_{t})$. Following standard practice [42], the reward is decomposed as:
$\mathrm{r}_{t}=\mathrm{r}_{t}^{\text{imitation}}+\mathrm{r}_{t}^{\text{regularization}}+\mathrm{r}_{t}^{\text{survival}}.$ (2)
Among these components, $\mathrm{r}_{t}^{\text{imitation}}$ encourages accurate imitation of the reference walking motion, which is generated using a closed-form ZMP (Zero Moment Point) solution [47]. $\mathrm{r}_{t}^{\text{regularization}}$ incorporates heuristics of ideal walking motion, penalizes joint torques, and promotes smooth actions to minimize unnecessary movements. A survival reward $\mathrm{r}_{t}^{\text{survival}}$ prevents early episode termination during training. Additional details are provided in Appendix 8.12.
4.3 Imitation Learning
Real-world data collection involves a human operator guiding the leader’s arms to teleoperate the follower’s arms while using a joystick and buttons on a handheld game controller to control body movements. During data collection, when the upper body tracks the position commands from the leader arms, ToddlerBot’s lower body employs a two-layer PD controller to actively maintain balance. The first layer is a Center of Mass (CoM) PD controller, which keeps the CoM close to the center of the support polygon. The second layer is a torso pitch PD controller, which uses IMU readings to ensure the torso remains upright. The first layer addresses CoM shifts caused by arm movements, while the second layer compensates when lifting heavy objects. With this setup, we can collect 60 trajectories in just 20 minutes for both bimanual and full-body manipulation tasks. The motor positions of the leader’s arms are recorded as the actions, while the motor positions of the follower robot, along with the RGB images captured from its camera, are recorded as observations. We train a diffusion policy [48] on this dataset, with further details provided in Appendix 8.13.
5 Experiments
Figure 3: Arm Span, Payload, and Trajectory Tracking. On the left, we show that with a torso dimension of $13\times 9\times 12\penalty 10000\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\penalty 10000\ \mathrm{cm}^{3}$, about 14 times the torso size. Additionally, ToddlerBot can lift weights up to $1484\penalty 10000\ \mathrm{g}$, which is 40% of its body weight ($3484\penalty 10000\ \mathrm{g}$). On the right, we present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile. Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
Capability: Arm Span, Payload, and Endurance. To evaluate ToddlerBot’s arm span, we teleoperate it to hug a large object using the compliant palm gripper while maintaining balance. We show that, ToddlerBot can grasp objects up to $27\times 24\times 31\penalty 10000\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance. ToddlerBot successfully lifts up to $1484\penalty 10000\ \mathrm{g}$, 40% of its total weight ($3484\penalty 10000\ \mathrm{g}$). To eliminate friction effects, we use a 3D-printed cup for the gripper to lock in securely. We incrementally add screws to the cup until ToddlerBot falls over. Results are shown on the left of Figure 3.
In the endurance test, ToddlerBot starts with a fully charged battery, running the walking RL policy while stepping in place. ToddlerBot achieved the longest streak of 19 minutes without falling. Over time, increased motor temperatures gradually pushed it outside the policy’s training distribution, leading to more frequent falls. ToddlerBot withstands up to 7 falls before breaking, but even then, repairs are quick—requiring only 21 minutes of 3D printing and 14 minutes of assembly, including removing the damaged part, installing the replacement, and performing zero-point calibration.
Capability: Push-ups and Pull-ups. To demonstrate expressive and dynamic motions, we program push-ups and pull-ups in our keyframe software and perform zero-shot sim-to-real transfer (Figure 4). For pull-ups, we use an AprilTag to help ToddlerBot locate the horizontal bar. Both tasks require strong limbs, balanced upper-lower body strength, and precise coordination, particularly when ToddlerBot transitions from a planking pose to standing after push-ups and when it releases the horizontal bar and lands after pull-ups. These open-loop transfers require only a single motion trajectory designed in simulation, highlighting the fidelity of our digital twin.
Figure 4: Experiment Results. We present four different tasks: push-up, pull-up, bimanual, and full-body manipulation, showing ToddlerBot’s capability in challenging loco-manipulation tasks.
Figure 5: Long-horizon Collaboration. In this task, two instances of ToddlerBot, Arya and Toddy, collaborate to clean up a toy session. (1) The task begins with a pink octopus on the table and a purple octopus on the ground. (2) Arya picks up the pink octopus from the table and places it in the wagon. (3) Arya walks to the wagon handle. (4) Arya grasps the handle while Toddy walks over. (5) Arya pushes the wagon toward the purple octopus. (6) Toddy reaches the pickup position. (7) Toddy kneels and picks up the purple octopus. (8) Finally, Arya and Toddy leave side by side.
ML-Compatibility: Omnidirectional Walking. To demonstrate ToddlerBot’s locomotion capabilities, we train RL walking policies to follow a square trajectory with a predefined velocity profile. The right part of Figure 3 presents the results, with real-world tracking data collected via motion capture. Due to RL policy limitations, both simulation and real-world tracking deviate from the command, primarily because the learned walking policy struggles with in-place rotation, causing translation offsets. However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer. Additionally, we report a position tracking error variance of $0.018\penalty 10000\ \mathrm{m}$, which demonstrates good repeatability (more details in Appendix 8.14).
ML-Compatibility: Vision-based Manipulation. We show ToddlerBot’s ability to perform bimanual manipulation and full-body manipulation by transferring octopus toys from a table and the ground to a wagon. Both tasks are trained with RGB-based diffusion policy [48] with 60 demonstrations. Across 20 test trials, we achieve a 90% success rate for bimanual manipulation and 75% for full-body manipulation. We leverage a combination of open-loop motions and closed-loop policies to enhance data collection efficiency. In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop. ToddlerBot’s onboard computing runs a 300300M parameter diffusion policy with about $100\penalty 10000\ \mathrm{ms}$ latency, enabling real-time operation. Results are shown in Figure 4.
ML-Compatibility: Skill Chaining We test ToddlerBot’s ability to combine loco-manipulation skills with wagon pushing. The results are shown in Appendix 8.15. To push the wagon, ToddlerBot first executes a diffusion policy to grasp the handle, while maintaining that pose, switches to the RL policy to walk forward. To enable walking while maintaining the grip, we sample the robot’s end pose from the grasping policy’s training data (60 demonstrations) during RL training.
Reproducibility: Hardware and Policies To demonstrate hardware reproducibility, we recruit a CS-major student with no prior hardware experience to assemble a second ToddlerBot using our open-source assembly manual and videos. The student independently completes the assembly in three days, including the time for 3D printing. The open-source community has also reported successful reproductions of ToddlerBot, and most of them were completed in one week (Appendix 8.16). For policy reproducibility, we run the manipulation policy trained on data collected with one instance on the other instance, achieving the same success rate of 90% across 20 trials. We also successfully transferred the RL walking policy between both robots. To further showcase the equivalent performance of both ToddlerBot instances, we have them collaborate on a long-horizon toy tidy-up session, as shown in Figure 5.
6 Conclusions
In conclusion, we demonstrate that ToddlerBot is ML-compatible, capable, and reproducible through a series of tests and loco-manipulation tasks. ToddlerBot extends beyond locomotion and supports full-body manipulation, character animation, human-robot interaction, and a range of ML applications, making it a versatile research platform. To bridge the sim-to-real gap, we develop a comprehensive sysID pipeline to create a high-fidelity digital twin, alongside a whole-body teleoperation system to collect loco-manipulation data. Being fully open-source, ToddlerBot empowers researchers to explore new research directions and fosters open collaboration across the community.
7 Limitations and Future Work
ToddlerBot’s performance in more agile tasks is constrained by the off-the-shelf motors’ max speed, max torque, and communication speed. Should there be stronger motor choices, the performance can still improve. Rather than achieving superhuman capabilities, ToddlerBot aligns more closely with average human performance in loco-manipulation tasks. In addition, the current actuation model does not consider motor temperature and is less accurate when close to the performance limit. The miniature scale limits its interaction with human-sized objects, though this does not hinder research if appropriately sized objects are used. Compared to metal shells, 3D-printed parts are more likely to break after the event of impact. Despite 3D printing being fast, repairs still require time.
To overcome these limitations, we plan to develop customized communication boards to increase control frequency and improve our actuation model to maximize usable envelope, especially when the motor heats up. We will refine ToddlerBot’s design to improve structural strength, as we believe that design has a greater impact than material choice. We also aim to improve sensing capabilities, including stereo vision for depth perception, additional IMUs for improved state estimation, and tactile sensors for richer manipulation feedback.
Acknowledgments
The authors would like to express their gratitude to Kaizhe Hu for assembling the second instance of ToddlerBot and assisting with keyframe animation and demo recording. We also extend our thanks to Huy Ha, Yen-Jen Wang, Pei Xu, and Yifan Hou for their insightful discussions on locomotion, and to Sirui Chen, Chen Wang, and Yunfan Jiang for valuable input on manipulation policy deployment. We are grateful to Albert Wu for his guidance on mathematical formulation and notation. Additionally, we thank João Pedro Araújo for his assistance with the motion capture system. Finally, we appreciate the helpful discussions from all members of TML and REALab. This work was supported by National Science Foundation NSF-FRR-2153854, NSF-2143601, NSF-2037101, NSF-2132519, Sloan Fellowship, Stanford Institute for Human-Centered Artificial Intelligence, and Stanford Wu Tsai Human Performance Alliance.